a bullet of mass 10g fired with an initial velocity of
500 ms-1 hits a 20 kg wooden block at rest  and gets embedded into it.
 
  1. calculate the velocity of the block after impact.
  2. how much energy is lost in the collision?

Grade:12

1 Answers

Susmita
5 years ago
Before collision mass of bullet m,speed v.After collision mass of bullet+block=m+M and speed V.
From conservation of linear momentum
mv=(m+M)V
Or,V=mv/(m+M)
m=10 g=10/1000 kg=0.01kg,M=20 kg,v=500m/s
So V=0.25 m/s
loss of kinetic enegy
=(1/2)[mv2-(m+M)V2]=1249.4 J
Please approve if you are helped
.